# GSPR: Multimodal Place Recognition Using 3D Gaussian Splatting for Autonomous Driving

### v1

@@ III-B3 Transformer Module
GSPR (ours) $\checkmark$ V+L 97.05 99.16 99.72 97.22 99.49 99.75 96.95 99.09 99.70
@@ III-C Two-step Training Strategy
We adopt a two-stage process to train the GSPR. First, we train explicit representations of autonomous driving scenes based on Multimodal Gaussian Splatting. Subsequently, the Global Descriptor Generator is trained for place recognition using the generated MGS scene representations.
@@
where $I_{\text{render}}$ is the image rendered by the 3D-GS renderer from the MGS scene representation, and $I_{\text{gt}}$ represents the ground-truth image. By evaluating the differences between the rendered image and the ground-truth image, we supervise the updates of Gaussian attributes and adjust the Gaussian distribution density. This process enables the iterative refinement of an accurate MGS scene representation of the real autonomous driving scenario.
@@ IV Experiments
IV-A Experimental Setup and Implementation Details
@@
We use nuScenes dataset [38] to evaluate our method, which consists of autonomous driving scenes collected from four different locations in Boston and Singapore: Boston Seaport (BS), SG-OneNorth (SON), SG-Queenstown (SQ), and SG-HollandVillage (SHV). The dataset provides multimodal data from 32-beam LiDAR and multi-view cameras, as well as rich 3D object annotations, facilitating scene reconstruction and allowing for comprehensive evaluation of place recognition methods. To obtain statistically significant results, we conduct experiments on the BS, SON, and SQ splits, which have sufficient loop closures and diverse situations.
We design the data preparation pipeline following previous works [16, 8], with some modifications. Specifically, we divide the data into the trainval set and the test set according to data collection time, and extract train query set and val query set from the trainval set, and test query set from the test set. We decrease the place density of the global map to 3 meters, and the place density of the query frames to 9 meters, thus widening the gap between viewpoints and raising the difficulty of place recognition. Additionally, both the database set and the training query set are utilized as query frames to ensure that the training set contains a sufficient number of triplets.
@@
For the Multimodal Gaussian Splatting module, we set the training iterations to 4000, the densification interval to 200, and stop densifying within 2000 iterations. This configuration is employed to reduce the training time of MGS and regulate the number of Gaussians. For the 3D graph convolution backbone, we set the number of neighbors $J=25$, the kernel support number $S=1$, and the sampling rate of the 3D graph max-pooling $r_{\text{pool}}=0.25$. For the transformer module, we set the positional embedding and the feature embedding dimension $d_{\text{pe}}=d_{\text{model}}=512$, the feed-forward dimension $d_{\text{ffn}}=1024$, and the number of heads $n_{\text{head}}=8$. For the NetVLAD module, we set the number of clusters $d_{\text{cluster}}=64$, and the output dimension $d_{\text{out}}=256$. An ADAM optimizer is used to train the network, while the initial learning rate is set to $1\times 10^{-7}$ and decays by a factor of 0.5 every 5 epochs. For the lazy triplet loss, we set the number of positive samples $k_{\text{pos}}=2$, the number of negative samples $k_{\text{neg}}=6$, and the margin $\beta=0.5$.
IV-B Evaluation for Place Recognition
@@
To validate the place recognition performance of GSPR in large-scale outdoor environments, we compare it with state-of-the-art baseline methods, including the visual-based methods NetVLAD [2] and SeqNet [37], the LiDAR-based methods OverlapTransformer [6], SeqNet-enhanced OverlapTransformer, and SeqOT [17], the radar-based method Autoplace [16], and the multimodal methods MinkLoc++ [7] and LCPR [8]. Among these, SeqNet [37], SeqNet-enhanced OverlapTransformer, SeqOT [17], and Autoplace [16] use sequential observations as inputs, compared to the other baselines only using one single frame for each retrieval. All experiments are conducted on a system with an Intel i7-14700KF CPU and an Nvidia RTX 4060Ti GPU.
We first train and test GSPR and the baseline methods on the largest BS split, and use the pretrained weights to perform zero-shot generalization experiments on the SON and SQ splits. We set the sequence length for all sequence-enhanced place recognition methods to 3 for evaluation fairness.
Following previous works [16, 18], we use average top 1 recall (AR@1), top 5 recall (AR@5), and top 10 recall (AR@10) as metrics to evaluate place recognition performance, with results presented in Tab. I. Our proposed GSPR outperforms all baselines on the BS, SON, and SQ splits, which include challenging conditions such as rain and nighttime. This demonstrates that our method effectively handles scenarios where unimodal approaches often fall short. Additionally, zero-shot generalization experiments on the SON and SQ splits indicate that our method achieves cross-scene and long-term generalization capability with limited training data.
IV-C Study on Maximum Sampling Range
We conduct an ablation study to investigate the impact of varying the maximum sampling range during voxel partitioning on GSPR’s place recognition performance. This study aims to evaluate the contribution of Gaussians distributed at different distances within the scene. The experimental results on the BS split are presented in Fig. 6. As can be observed, the best place recognition performance occurs when the maximum sampling distance is at least 40 meters. Notably, the AR@1 does not significantly improve with sampling range increase after 40 meters. A possible reason is that each Gaussian scene is initialized from LiDAR data where points at greater distances are more sparse, leading to less distinct spatio-temporal features to boost the recognition performance.
TABLE II: The impact of different voxel numbers
Voxel Number AR@1 AR@5 AR@10 AR@20 Inference Time [ms]
1024 72.09 91.58 95.93 97.76 4.32
2048 88.22 97.62 98.88 99.72 4.52
4096 96.77 98.60 99.16 99.16 11.76
8192 97.05 99.16 99.72 99.72 41.02
IV-D Study on Number of Voxels
In our experimental setup, each scene consists of approximately 100,000 Gaussians. To organize the unordered Gaussian scene and mitigate computational load, we downsample it into a set of $N$ voxels using voxel partitioning. Inspired by the LiDAR downsampling algorithm in PointNetVLAD [4], we adaptively adjust voxel size to transform all Gaussian scenes into $N$ voxels. A larger $N$ preserves fine details but incurs higher computational costs, while a smaller $N$ increases inference speed at the cost of information loss. To assess the impact of voxel count $N$ on place recognition performance, we conduct experiments on the BS split. As shown in Tab. II, setting $N$ to 4096 achieves a balance between performance and inference speed, as very low voxel counts result in significant information loss, while excessively high counts lead to impractical computational demands.
TABLE III: Ablation study of input features on the BS split
@@ IV-D Study on Number of Voxels
IV-E Ablation Study on Input Features
A Gaussian $g=[\mu^{\scriptscriptstyle\text{G}},s,q,sh,\alpha]\in\mathbb{R}^{59}$ is composed of different parts of features, including position $\mu^{\scriptscriptstyle\text{G}}\in\mathbb{R}^{3}$, scale $s\in\mathbb{R}^{3}$, rotation $q\in\mathbb{R}^{4}$, SH coefficients $sh\in\mathbb{R}^{48}$, and opacity $\alpha\in\mathbb{R}^{1}$. In GSPR, position is fed into the transformer module as positional encoding, while scale, rotation, SH coefficients, and opacity are input features of the graph convolution backbone network. We conduct an ablation study on input features using the BS split to examine the impact of different components of the Gaussian on place recognition performance. The results are shown in Tab. III. As can be seen, in addition to the SH coefficients, the Gaussian features of position, scale, and opacity provide valuable information for place recognition, while the rotation feature contributes relatively less. This suggests that ”where it is” is more expressive than ”which direction it is heading” for Gaussian-based place description, as the former corresponds more directly to the explicit spatial structure of the places.
@@ V Conclusion
In this paper, we present GSPR, a novel multimodal place recognition network based on 3D-GS. Our method proposes Multimodal Gaussian Splatting to harmonize continuous observations, including multi-view RGB images and LiDAR point clouds, creating a spatio-temporal unified MGS scene representation tailored for autonomous driving scenarios. To manage the unordered Gaussians, we implement voxel downsampling for effective organization. We further propose using 3D graph convolution networks and transformer module to exploit local and global spatio-temporal features from Gaussian graphs, generating discriminative descriptors for place recognition. The experimental results suggest that our method outperforms state-of-the-art baselines, including unimodal, multimodal, and sequence-enhanced methods. It demonstrates the advantages of the 3D-GS-based multimodal fusion approach in handling challenging place recognition tasks with notable generalization ability.

### v2

@@ III-B Global Descriptor Generator
GSPR-L (ours) $\checkmark$ V+L 96.05 99.16 99.30 93.94 97.98 99.24 94.21 98.17 99.70
GSPR (ours) $\checkmark$ V+L 98.74 99.44 99.44 98.99 99.75 100.00 99.09 99.70 99.70
TABLE II: Comparision of place recognition performance on the KITTI and KITTI-360 datasets
Methods Sequence1 Modality2 KITTI KITTI-360 Generalization
AR@1 AR@5 AR@10 AR@1 AR@5 AR@10 AR@1 AR@5 AR@10
MinkLoc++ [7] $\times$ V+L 95.76 99.15 99.72 95.89 99.03 99.27 96.37 99.27 99.52
SeqOT [18] $\checkmark$ L 97.46 99.15 99.72 98.07 99.27 99.40 98.31 99.52 99.88
GSPR-L (ours) $\checkmark$ V+L 98.31 99.72 99.72 98.55 99.40 99.40 97.34 99.15 99.27
GSPR (ours) $\checkmark$ V+L 99.44 99.72 100.00 99.15 99.64 99.64 98.91 99.64 99.64
@@ III-C Two-step Training Strategy
We adopt a two-stage process to train the GSPR. Firstly, we train explicit representations of autonomous driving scenes based on Multimodal Gaussian Splatting. Subsequently, the Global Descriptor Generator is trained for place recognition using the generated MGS scene representations.
@@
where $I_{\text{render}}$ is the image rendered by the 3D-GS renderer from the MGS scene representation, and $I_{\text{gt}}$ represents the ground-truth image. We use $\mathcal{L}_{\text{MGS}}$ to supervise the iterative refinement of MGS scene representations to accurately reconstruct the autonomous driving scenarios.
@@ IV Experiments
IV-A Experimental Setup
We use three datasets, nuScenes [36], KITTI [37], and KITTI-360 [38], to evaluate the place recognition accuracy and generalization performance of our proposed GSPR.
@@ IV-A Experimental Setup
nuScenes. It includes autonomous driving scenes collected from four different locations: Boston Seaport (BS), SG-OneNorth (SON), SG-Queenstown (SQ), and SG-HollandVillage (SHV). It provides multimodal data from 32-beam LiDAR and multi-view cameras. To obtain statistically significant results, we conduct experiments on the BS, SON, and SQ splits, which have sufficient loop closures and diverse situations. Our data preparation pipeline mainly follows [17, 8]. Inspired by [39], we construct a sparse scan map by downsampling the database set at 3 meter intervals and the test set at 9 meter intervals, to evaluate recognition accuracy under large viewpoint differences. To construct a sequence, we use the current observation, along with the previous and next observations that are temporally adjacent to it.
KITTI. It is a standard autonomous driving dataset that includes various urban scenarios and traffic conditions. It provides multimodal data, including front-view stereo cameras and the Velodyne HDL-64E LiDAR, with GNSS-based ground-truth poses. We select Sequence 02 for training and Sequence 00 for testing, and perform data partitioning in the same manner as for the nuScenes dataset.
KITTI-360. It is a larger multimodal dataset compared to KITTI, with a similar sensor configuration. We select 2013-05-28-drive-$0$ for training and 2013-05-28-drive-0002 for testing. Additionally, we transfer the weights trained on the KITTI dataset to KITTI-360 to evaluate the cross-dataset generalization ability of our proposed method. Notably, the field of view of the cameras in KITTI and KITTI-360 is considerably smaller than that of nuScenes, making the panoramic reconstruction results unavailable. We demonstrate the robustness of GSPR to sensor configuration through experiments on KITTI and KITTI-360.
IV-B GSPR Implementation Details
@@
For the Gaussian Optimization module, we set the training iteration to 400, which significantly accelerates the training and inference of the network. This setting sacrifices some rendering quality, but is sufficient to reconstruct the dense and uniform Gaussian scene representations. For the 3D graph convolution backbone, we set the number of neighbors $J=25$, the kernel support number $S=1$, and the sampling rate of the 3D graph max-pooling $r_{\text{pool}}=0.25$. For the transformer module, we set the positional embedding and the feature embedding dimension $d_{\text{pe}}=d_{\text{model}}=512$, the feed-forward dimension $d_{\text{ffn}}=1024$, and the number of heads $n_{\text{head}}=8$. For the NetVLAD module, we set the number of clusters $d_{\text{cluster}}=64$, and the descriptor dimension $d_{\text{out}}=256$. An ADAM optimizer is used to train the network, while the initial learning rate is set to $1\times 10^{-5}$ and decays by a factor of 0.5 every 5 epochs. For the lazy triplet loss, we set the number of positive samples $k_{\text{pos}}=2$, the number of negative samples $k_{\text{neg}}=6$, and the margin $\beta=0.5$.
In addition, we set the number of input voxels in GSPR to $N=4096/8192$ during training/inference respectively. We also design a lightweight version, GSPR-L, with only half the number of input voxels in GSPR for inference. All experiments are conducted on a system with an Intel i7-14700KF CPU and an Nvidia RTX 4060Ti GPU.
IV-C Evaluation for Place Recognition
@@
To validate the place recognition performance of GSPR in large-scale outdoor environments, we compare it with state-of-the-art baseline methods, including the visual-based methods AnyLoc [3] and SeqNet [35], the LiDAR-based methods OverlapTransformer [6], and SeqOT [18], the radar-based method Autoplace [17], and the multimodal methods MinkLoc++ [7] and LCPR [8]. Among these, SeqNet [35], SeqOT [18], and Autoplace [17] use sequential observations as inputs, compared to the other baselines only using one single frame for each retrieval. We try to reproduce the baselines using their open source code. During the mixed masking process of GSPR, we use ground-truth 3D annotations for training on nuScenes. For inference on nuScenes and the entire deployment on KITTI and KITTI-360, we use annotations generated by PointPillars [40]. We set the sequence length for all sequence-enhanced place recognition methods to 3 for fairness.
Following previous works [17, 19], we use average top 1 recall (AR@1), top 5 recall (AR@5), and top 10 recall (AR@10) as metrics to evaluate the place recognition performance. The results on the nuScenes dataset are shown in Tab. I. In addressing challenging scenarios that include rain and nighttime conditions, our proposed GSPR holds the best recognition accuracy on all metrics, while GSPR-L strikes a balance between inference speed (approximately one-third of the GSPR runtime for global descriptor generation) and recognition accuracy. This demonstrates that our method effectively handles scenarios where unimodal approaches fail, and achieves good recognition accuracy under large viewpoint differences.
The experimental results on the KITTI and KITTI-360 datasets are shown in Tab. II. Our method presents good place recognition accuracy on both KITTI and KITTI-360 datasets, while showing strong generalization performance in cross-dataset scenarios. Furthermore, the experimental results on KITTI and KITTI-360 also demonstrate the solid robustness of GSPR in the case where panoramic reconstruction results are not available. When the 3D-GS reconstruction is supervised using only a front-view stereo camera, GSPR still maintains the best recognition accuracy.
TABLE III: Ablation study of improvement strategies
LiDAR Initialization Spherical Dome Static Mask Dynamic Mask AR@1 AR@5 AR@10
12.81 28.77 39.02
91.85 98.04 98.88
92.13 98.46 99.16
93.67 98.88 99.02
96.05 99.16 99.30
IV-D Ablation Studies
Improvement strategies. We ablate the improvement strategies of our MGS module in generating Gaussian scenes tailored for place recognition. The experimental results of GSPR-L in Tab. III show that each improvement strategy of MGS has a positive effect on place recognition performance. In particular, using the Gaussian scenes generated by the vanilla 3D-GS as input results in a relatively low recognition accuracy (the first row of Tab. III). This is probably due to the difficulty of SfM in producing reliable sparse reconstruction results on the nuScenes dataset, resulting in poor Gaussian initialization and suboptimal scene reconstruction.
Input features. A Gaussian $g=[\mu^{\scriptscriptstyle\text{G}},s,q,sh,\alpha]\in\mathbb{R}^{59}$ is composed of different parts of features, including position $\mu^{\scriptscriptstyle\text{G}}\in\mathbb{R}^{3}$, scale $s\in\mathbb{R}^{3}$, rotation $q\in\mathbb{R}^{4}$, SH coefficients $sh\in\mathbb{R}^{48}$, and opacity $\alpha\in\mathbb{R}^{1}$. We ablate these input features using the BS split to assess their impact on recognition performance. The results of GSPR-L shown in Tab. IV indicate that, in addition to the SH coefficients, position, scale, and opacity are crucial for place recognition, while the rotation feature contributes less. This suggests that “where it is” is more expressive than “which direction it is heading” for Gaussian-based place description, as the former corresponds more directly to the explicit spatial structure of the places.
TABLE IV: Ablation study of input features on the BS split
@@ IV-D Ablation Studies
Maximum sampling ranges. We further explore the impact of varying the maximum sampling range during voxel partitioning on GSPR’s recognition performance, focusing on the contribution of Gaussians distributed at different distances within the scene. As shown in Fig. 6, the best place recognition performance occurs when the maximum sampling distance is at least 40 meters. Notably, the AR@1 does not significantly increase with sampling range increase after 40 meters. A possible reason is that each Gaussian scene is initialized from LiDAR data where points at greater distances are more sparse, leading to less distinct spatio-temporal features to boost the recognition performance.
@@ V Conclusion
In this paper, we present GSPR, a novel multimodal place recognition network based on 3D-GS. Our method proposes Multimodal Gaussian Splatting to harmonize multi-view RGB images and LiDAR point clouds into a unified spatio-temporal MGS scene representation tailored for place recognition. To manage the unordered Gaussians, we apply voxel downsampling for efficient data organization. We further propose using 3D graph convolution networks and transformer module to exploit local and global spatio-temporal features from Gaussian graphs, generating discriminative global descriptors. Experimental results indicate that our method outperforms state-of-the-art baselines, demonstrating the advantages of the 3D-GS-based multimodal fusion approach for challenging place recognition tasks.